# Is Repeated Bayesian Interim Analysis Consequence-Free?

### v1

@@ Abstract
Interim analyses are commonly used in clinical trials to enable early stopping for efficacy, futility, or safety. While their impact on frequentist operating characteristics is well studied and broadly understood, the effect of repeated Bayesian interim analyses—when conducted without appropriate multiplicity adjustment— remains an area of active debate. In this article, we provide both theoretical justification and numerical evidence illustrating how such analyses affect key inferential properties, including bias, mean squared error, the coverage probability of posterior credible intervals, false discovery rate, familywise error rate, and power. Our findings demonstrate that Bayesian interim analyses can significantly alter a trial’s operating characteristics, even when the prior used for Bayesian inference is correctly specified and aligned with the data-generating process. Extensive simulation studies, covering a variety of endpoints, trial designs (single-arm and two-arm randomized controlled trials), and scenarios with both correctly specified and misspecified priors, support theoretical insights. Collectively, these results underscore the necessity of appropriate adjustment, thoughtful prior specification, and comprehensive evaluation to ensure valid and reliable inference in Bayesian adaptive trial designs.
@@ 1 Introduction
In clinical trials, interim analyses are often conducted to assess trial progress and inform decisions about whether to continue or stop the trial early. A trial may be stopped early if interim results cross predefined thresholds, such as showing compelling efficacy, indicating futility due to lack of benefit, or revealing unacceptable toxicity that compromises participant safety. Two primary types of decision rules are used in interim analyses and stopping: frequentist and Bayesian.
@@
In the frequentist approach, the impact of interim analyses on the operating characteristics of clinical trials—particularly the familywise error rate (FWER)—has been extensively studied and is generally understood. Within this framework, parameters are treated as fixed but unknown constants. Repeated interim looks introduce a multiplicity problem, leading to inflation of the FWER. To address this, various methods have been developed to adjust significance levels or decision rules at each interim analysis, thereby ensuring that the overall FWER remains controlled at the desired level. The Pocock method (Pocock, 1977) assigns equally stringent significance levels to all interim analyses, making early stopping easier but potentially sacrificing efficiency and power. In contrast, the O’Brien-Fleming method (1979) applies more stringent stopping criteria early in the trial and gradually relaxes them as the trial progresses, offering a more conservative approach. Alpha spending functions (DeMets and Lan, 1983) provide additional flexibility by allocating the total alpha level across multiple analyses and dynamically adjusting critical values based on how much alpha has already been ‘spent’.
@@
In contrast, the Bayesian framework—which treats parameters as random variables with associated probability distributions—has been less extensively studied in the context of interim analyses, and there is no clear consensus on their impact on trial operating characteristics. Berry (1987) argued that Bayesian interim analyses are not subject to multiplicity issues, as they adhere to the likelihood principle, allowing analyses to be conducted at any point during the trial without requiring adjustments. Harrell (2017) echoed this perspective, advocating the use of Bayesian posterior probabilities to make go/no-go decisions as frequently as needed. However, this view has not gone unchallenged. Jennison and Turnbull (1990) cautioned that Bayesian methods may exhibit poor frequentist properties, including an inflated FWER when multiple interim looks are performed.
Beyond FWER control, less attention has been given to other statistical properties, such as bias, mean square error (MSE), and interval coverage in treatment effect estimation. In the frequentist framework, Armitage (1958) pointed out that multiple interim looks can impact the maximum likelihood estimator and the coverage probability of the confidence interval. Tsiatis, Rosner, and Mehta (1984) described a method of constructing confidence intervals with exact coverage probabilities for the mean of a normal distribution. In the Bayesian framework, Rosenbaum and Rubin (1984) conducted simulation studies to examine the coverage probability of the posterior credible interval under a flat user-specified prior, when the data-generating prior was a normal distribution with varying precision. They found that coverage probabilities deviated from nominal value when repeated Bayesian interim analyses were applied, particularly when the precision of the data-generating prior increased. More recently, Bunouf (2023) proposed the Beta-J prior for single-arm binomial trials, demonstrating improved frequentist properties, such as reduced bias and enhanced interval coverage, compared to standard conjugate Beta priors.
Despite these findings, most research has prioritized FWER control over evaluating other critical inferential quantities, such as bias, MSE, and interval coverage. Recent work by Robertson et al. (2023) and Bunouf (2023) has highlighted this gap, calling for more comprehensive assessments of how early trial termination affects treatment effect estimates. The U.S. Food and Drug Administration has also acknowledged this concern, noting in recent guidance (FDA, 2019) that, compared to the FWER inflation, “biased estimation in adaptive design is currently a less well-studied phenomenon.”
In this article, we aim to address this critical knowledge gap by providing both theoretical and numerical examinations of how repeated Bayesian interim analyses affect key statistical properties of clinical trial designs. Specifically, we assess their impact on bias, MSE, coverage probability of the posterior credible intervals, false discovery rate (FDR), FWER, and power. Our analysis explicitly distinguishes between the data-generating prior and the user-specified prior. We show that even when the two priors are identical, repeated Bayesian interim analyses can systematically distort some key inferential quantities. These distortions become even more pronounced when the user-specified prior is misspecified. To support our theoretical findings, we conduct extensive simulation studies across a wide range of settings, including different types of endpoints (binary, continuous, and time-to-event) and both single-arm and two-arm randomized controlled trials (RCTs), highlighting the generality and practical significance of our results.
@@
The remainder of this article is organized as follows. In Section 2, we present Bayesian probability models for binary, normal, and time-to-event outcomes, along with a theoretical assessment of the impact of repeated Bayesian interim analyses on the operating characteristics of clinical trials. Section 3 evaluates these effects through extensive simulation studies. Finally, we offer concluding remarks in Section 4.
@@ 2.1 Probability Models
Consider a superiority trial that aims to demonstrate the superiority of an experimental treatment over a control, with interim analyses that may lead to early termination if there is overwhelming evidence of superiority. Let $X_{i}$ denote the efficacy endpoint for the $i$th patient receiving the experimental treatment, where $X_{i}$ follows a probability distribution indexed by the parameter $\theta$, representing the treatment effect of interest, possibly along with other nuisance parameters. Without loss of generality, we assume that larger values of $\theta$ indicate higher efficacy. The hypotheses of interest are formulated as:
$H_{0}:\theta-\theta_{0}\leq\delta\quad\mathrm{vs.}\quad H_{a}:\theta-\theta_{0}>\delta,$
where $\theta_{0}$ denotes the treatment effect for the control, and $\delta$ is an equivalence margin, which is often set at 0 in superiority trials, and a negative value for non-inferiority trials. In single-arm trials, $\theta_{0}$ is assumed to be a known constant, while in two-arm RCTs, $\theta_{0}$ is treated as an unknown parameter. While we adopt this formulation for clarity of exposition, alternative hypothesis structures, such as those based on hazard ratios for time-to-event endpoints in RCTs, are also possible and are discussed later.
In the Bayesian framework, the parameter $\theta$ is treated as a random variable with an associated probability distribution. For our analysis, it is essential to distinguish between two types of priors for $\theta$: the data-generating prior and the user-specified prior. The data-generating prior distribution of $\theta$ represents the true underlying distribution that generates $\theta$, which in turn generates the observed data $X_{i}$. In practice, the data-generating prior is typically unknown. The user-specified prior, on the other hand, refers to the prior assumed by users for Bayesian inference. It can be constructed based on prior knowledge, historical data, or chosen to be noninformative or vague to limit its impact on the posterior. In nearly all applications, the user-specified prior is misspecified in that it differs from the true data-generating prior.
Next, we present commonly used Bayesian models for binary, normal, and time-to-event endpoints, along with their associated prior and posterior distributions. For binary and normal outcomes in RCTs, patients in the treatment and control arms are independent, allowing the same model to be applied separately to each arm. As a result, we do not describe separate models for single-arm trials and RCTs in these cases. For time-to-event endpoints, considering the popularity of using the Cox proportional hazards model (Cox, 1972) in RCTs, we present separate modeling approaches for single-arm trials and RCTs. These models are not required for the theoretical development presented in the next section, but they facilitate the understanding of theoretical results and serve as the foundation for the simulation study.
@@
$X_{i}\sim\mbox{Bernoulli}(\theta),$ (1)
@@
$\nu_{n}=\frac{n\bar{X}_{n}+\kappa\mu}{\kappa+n}$
$\tau_{n}=\big{\{}\nu\sigma_{0}^{2}+\sum(X_{i}-\bar{X}_{n})^{2}+\frac{\kappa n}{\kappa+n}(\bar{X}_{n}-\mu)^{2}\big{\}}/(\nu+n)(\kappa+n).$
@@
$\mu_{n}=\frac{\frac{1}{\sigma_{0}^{2}}\mu+\frac{n}{\sigma^{2}}\bar{X}_{n}}{\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}}\ \ \ \ \ \ \mbox{and}\ \ \ \ \ \ \frac{1}{\sigma_{n}^{2}}=\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}.$ (3)
@@
$X_{i}\sim\mbox{Weibull}(\theta,\kappa),$ (4)
with probability density function
$f(X_{i}|\theta,\kappa)=(ln2)\kappa\theta^{-\kappa}{X_{i}}^{\kappa-1}\mbox{exp}\big{\{}-(ln2)\theta^{-\kappa}{X_{i}}^{\kappa}\big{\}},$ (5)
and survival function
$S(X_{i}|\theta,\kappa)=\mbox{exp}\big{\{}-ln(2)\theta^{-\kappa}X_{i}^{\kappa}\big{\}}.$ (6)
This parameterization expresses the Weibull distribution in terms of the median $\theta$, which often has a more interpretable clinical interpretation, and the shape parameter $\kappa$. At any interim decision time, let $t_{i}$ denote the observed time to event or censoring time for the $i$th patient, and let $\zeta_{i}=I(t_{i}=X_{i})$ denote the censoring indicator. The likelihood for patient $i$, based on data $(t_{i},\zeta_{i})$, is given by $L\big{(}(t_{i},\zeta_{i})|\theta,\kappa\big{)}=f(t_{i}|\theta,\kappa)^{\zeta_{i}}S(t_{i}|\theta,\kappa)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big{(}(t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big{)}$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa)=\prod_{i=1}^{n}L\big{(}(t_{i},\zeta_{i})|\theta,\kappa\big{)}$.
We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$ and $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$ for the median $\theta$ and shape parameter $\kappa$, respectively. Given interim data ${\cal D}_{n}$, the joint posterior distribution of $(\theta,\kappa)$ under the user-specified prior is $\pi(\theta,\kappa|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)L({\cal D}_{n}|\theta,\kappa)$, which can be sampled using the Gibbs sampler (Geman and Geman, 1984; Gelfand and Smith, 1990).
@@
$h(t|\theta,\kappa,\beta)=h_{0}(t|\theta,\kappa)\mbox{exp}(\beta Z),$ (7)
where $Z$ is the treatment indicator ($Z=1$ for the experimental treatment and $Z=0$ for the control), and $\mbox{exp}(\beta)$ represents the hazard ratio (HR) comparing the experimental treatment to the control. We consider the following hypotheses test: $H_{0}:\mbox{exp}(\beta)>\rho$ versus $H_{a}:\mbox{exp}(\beta)<\rho$, where $\rho$ is a predefined upper bound for HR used to define superiority.
We specify the baseline hazard function $h_{0}(t|\theta,\kappa)=(ln2))\theta^{-\kappa}\kappa t^{\kappa-1}$, corresponding to a Weibull distribution parameterized by the median $\theta$ and shape parameter $\kappa$. The resulting probability density function and survival function are
$f(t|\theta,\kappa,\beta)=(ln2)\kappa\theta^{-\kappa}{t}^{\kappa-1}\mbox{exp}(\beta Z)\mbox{exp}(-(ln2)\theta^{-\kappa}\mbox{exp}(\beta Z){t}^{\kappa}),$ (8)
and
$S(t|\theta,\kappa,\beta)=\mbox{exp}\big{(}-ln(2)\theta^{-\kappa}\mbox{exp}(\beta Z)t^{\kappa}\big{)}.$ (9)
The likelihood for patient $i$ with data $(t_{i},\zeta_{i})$ is $L\big{(}(t_{i},\zeta_{i})|\theta,\kappa,\beta\big{)}=f(t_{i}|\theta,\kappa,\beta)^{\zeta_{i}}S(t_{i}|\theta,\kappa,\beta)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big{(}(t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big{)}$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa,\beta)=\prod_{i=1}^{n}L\big{(}(t_{i},\zeta_{i})|\theta,\kappa,\beta\big{)}$. We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$, $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$, and $\pi(\beta)=\mbox{Normal}(\xi,\tau^{2})$ for the parameters $\theta$, $\kappa$, and $\beta$, respectively. The joint posterior distribution of $(\theta,\kappa,\beta)$ under the user-specified prior is $\pi(\theta,\kappa,\beta|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)\pi(\beta)L({\cal D}_{n}|\theta,\kappa,\beta)$, which can be sampled using the Gibbs sampler.

### v2

@@ Abstract
Interim analyses are vital in clinical trials for early decision-making. While frequentist implications are well-established, the consequences of repeated Bayesian interim monitoring for efficacy, specifically regarding multiplicity, remain contentious. This article provides theoretical justification and numerical evidence evaluating the impact of such designs on bias, mean squared error (MSE), credible interval coverage, false discovery rate (FDR), and average Type I error (ATIE). Our findings show that when the inferential prior matches the data-generating prior, sequential efficacy stopping does not bias the posterior mean or degrade credible interval coverage. However, even under this “matched” condition, the FDR, ATIE, and MSE are significantly altered. In the more practically relevant scenario where the inferential and data-generating priors differ, all aforementioned operating characteristics, including estimation bias and coverage, are substantially impacted. These results reconcile long-standing conflicting arguments regarding Bayesian multiplicity. We demonstrate that while some Bayesian properties are invariant to sequential looks, others are not. Our work underscores the necessity of thoughtful prior specification and comprehensive evaluation of frequentist-Bayesian operating characteristics to ensure reliable inference in adaptive trial designs.
@@ 1 Introduction
In clinical trials, interim analyses are often conducted to assess trial progress and inform decisions about whether to continue or stop the trial early. A trial may be stopped early if interim results cross predefined thresholds, such as showing compelling efficacy, or indicating futility due to lack of benefit. In this article, we focus on interim analyses that may lead to early stopping and claim of efficacy, refer to as superiority stopping. Two primary types of decision rules are used in interim analyses for superiority stopping: frequentist and Bayesian.
@@
In the frequentist approach, the impact of interim analyses on the operating characteristics of clinical trials, particularly the familywise error rate (FWER), has been extensively studied and is generally understood. Within this framework, parameters are treated as fixed but unknown constants. Repeated interim looks with superiority stopping introduce a multiplicity problem, leading to inflation of the FWER. To address this, various methods have been developed to adjust significance levels or decision rules at each interim analysis, thereby ensuring that the overall FWER remains controlled at the desired level. The Pocock method \@BBOPcitep\@BAP\@BBN(Pocock77)\@BBCP assigns equally stringent significance levels to all interim analyses, making early stopping easier but potentially sacrificing efficiency and power. In contrast, the O’Brien-Fleming method \@BBOPcitep\@BAP\@BBN(OBrien79)\@BBCP applies more stringent stopping criteria early in the trial and gradually relaxes them as the trial progresses, offering a more conservative approach. Alpha spending functions \@BBOPcitep\@BAP\@BBN(Lan83)\@BBCP provide additional flexibility by allocating the total alpha level across multiple analyses and dynamically adjusting critical values based on how much alpha has already been ‘spent’.
@@
In contrast, the Bayesian framework, which treats parameters as random variables with associated probability distributions, has been less thoroughly examined in the context of interim analyses, and no clear consensus exists regarding their effects on trial operating characteristics. \@BBOPcitet\@BAP\@BBNBerry87\@BBCP argued that Bayesian interim analyses with superiority stopping are not subject to multiplicity concerns because they adhere to the likelihood principle, allowing analyses to be conducted at any point during the trial without adjustment. \@BBOPcitet\@BAP\@BBNHarrell17\@BBCP echoed this perspective, advocating the use of Bayesian posterior probabilities to make go/no-go decisions as frequently as needed.
However, this view has been met with skepticism. \@BBOPcitet\@BAP\@BBNArmitage58\@BBCP noted that multiple interim looks can affect the maximum likelihood estimator and the coverage probability of confidence intervals. \@BBOPcitet\@BAP\@BBNJennison90\@BBCP further cautioned that Bayesian methods may exhibit poor frequentist properties, including inflated FWER when multiple interim analyses are conducted. \@BBOPcitet\@BAP\@BBNRosenbaum84\@BBCP conducted simulation studies showing that the coverage probability of posterior credible intervals can be distorted when commonly used flat, non-informative priors are employed for inference. Most critiques in the literature focus primarily on frequentist quantities, particularly FWER, raising questions about their relevance and about the appropriateness of using frequentist metrics to assess Bayesian inference. The U.S. Food and Drug Administration has also acknowledged this issue, stating in recent guidance \@BBOPcitep\@BAP\@BBN(FDA19)\@BBCP that, relative to FWER inflation, “biased estimation in adaptive design is currently a less well-studied phenomenon.”
In this article, we aim to fill this critical knowledge gap by providing both theoretical and numerical investigations into how repeated Bayesian interim analyses affect the statistical properties of clinical trial designs. Specifically, we examine their impact on the bias and mean squared error (MSE) of the posterior mean, the coverage probability of posterior credible intervals, and error metrics such as the false discovery rate (FDR) and average type I error (ATIE)—quantities that, while often viewed through a frequentist lens, are evaluated here within a fully Bayesian framework. Our analysis explicitly distinguishes between the data-generating prior and the user-specified prior, accounting for the probability distribution of unknown parameters to maintain a strictly Bayesian perspective.
While some argue that Bayesian interim monitoring allows for continuous looks without adjustment, we show this claim holds only for specific quantities, such as the posterior mean and credible interval coverage. More importantly, it relies on the assumption that the user-specified (i.e., analyst’s) prior perfectly matches the true data-generating prior. This “matched-prior” assumption is rarely met in practice and has been largely unacknowledged in seminal studies, including \@BBOPcitet\@BAP\@BBNBerry87\@BBCP. By identifying these boundary conditions, our findings reconcile the conflicting arguments regarding Bayesian multiplicity that have long persisted in the literature.
Moreover, we demonstrate that even when the two priors coincide, repeated Bayesian interim analyses can systematically distort other key statistical properties, including the FDR and ATIE. Crucially, when the user-specified prior deviates from the data-generating prior, the typical scenario in practice, all evaluated metrics are adversely affected.
@@
The remainder of this article is organized as follows. In Section 2, we present Bayesian probability models for binary, normal, and time-to-event outcomes, along with a theoretical assessment of the impact of repeated Bayesian interim analyses on the operating characteristics of clinical trials. Section 3 evaluates these effects through extensive simulation studies. We offer concluding remarks in Section 4.
@@ 2.1 Probability Models
Consider a superiority trial that aims to demonstrate the superiority of an experimental treatment over a control, with interim analyses that may lead to early termination if there is overwhelming evidence of superiority. Let $X_{i}$ denote the efficacy endpoint for the $i$th patient receiving the experimental treatment, where $X_{i}$ follows a probability distribution indexed by the parameter $\theta$, representing the treatment effect of interest, possibly along with other nuisance parameters. Without loss of generality, we assume that larger values of $\theta$ indicate higher efficacy.
Next, we present commonly used Bayesian models for binary, normal, and time-to-event endpoints, along with their associated prior and posterior distributions. For binary and normal outcomes in randomized controlled trials (RCTs), patients in the treatment and control arms are independent, allowing the same model to be applied separately to each arm. As a result, we do not describe separate models for single-arm trials and RCTs in these cases. For time-to-event endpoints, considering the popularity of using the Cox proportional hazards model \@BBOPcitep\@BAP\@BBN(Cox72)\@BBCP in RCTs, we present separate modeling approaches for single-arm trials and RCTs. These models are not required for the theoretical development presented in the next section, but they facilitate the understanding of theoretical results and serve as the foundation for the simulation study.
@@
$X_{i}\sim\mbox{Bernoulli}(\theta),$
@@
$\nu_{n}=\frac{n\bar{X}_{n}+\kappa\mu}{\kappa+n},\quad\tau_{n}=\big\{\nu\sigma_{0}^{2}+\sum(X_{i}-\bar{X}_{n})^{2}+\frac{\kappa n}{\kappa+n}(\bar{X}_{n}-\mu)^{2}\big\}/(\nu+n)(\kappa+n).$
@@
$\mu_{n}=\frac{\frac{1}{\sigma_{0}^{2}}\mu+\frac{n}{\sigma^{2}}\bar{X}_{n}}{\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}}\ \ \ \ \ \ \mbox{and}\ \ \ \ \ \ \frac{1}{\sigma_{n}^{2}}=\frac{1}{\sigma_{0}^{2}}+\frac{n}{\sigma^{2}}.$
@@
$X_{i}\sim\mbox{Weibull}(\theta,\kappa),$
with probability density function $f(X_{i}|\theta,\kappa)=(ln2)\kappa\theta^{-\kappa}{X_{i}}^{\kappa-1}\mbox{exp}\big\{-(ln2)\theta^{-\kappa}{X_{i}}^{\kappa}\big\},$ and survival function $S(X_{i}|\theta,\kappa)=\mbox{exp}\big\{-ln(2)\theta^{-\kappa}X_{i}^{\kappa}\big\}$. This parameterization expresses the Weibull distribution in terms of the median $\theta$, which often has a more interpretable clinical interpretation, and the shape parameter $\kappa$. At any interim decision time, let $t_{i}$ denote the observed time to event or censoring time for the $i$th patient, and let $\zeta_{i}=I(t_{i}=X_{i})$ denote the censoring indicator. The likelihood for patient $i$, based on data $(t_{i},\zeta_{i})$, is given by $L\big((t_{i},\zeta_{i})|\theta,\kappa\big)=f(t_{i}|\theta,\kappa)^{\zeta_{i}}S(t_{i}|\theta,\kappa)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big((t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big)$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa)=\prod_{i=1}^{n}L\big((t_{i},\zeta_{i})|\theta,\kappa\big)$.
We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$ and $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$ for the median $\theta$ and shape parameter $\kappa$, respectively. Given interim data ${\cal D}_{n}$, the joint posterior distribution of $(\theta,\kappa)$ under the user-specified prior is $\pi(\theta,\kappa|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)L({\cal D}_{n}|\theta,\kappa)$, which can be sampled using the Gibbs sampler \@BBOPcitep\@BAP\@BBN(Geman84)\@BBCP
@@
$h(t|\theta,\kappa,\beta)=h_{0}(t|\theta,\kappa)\mbox{exp}(\beta Z),$
where $Z$ is the treatment indicator ($Z=1$ for the experimental treatment and $Z=0$ for the control), and $\mbox{exp}(\beta)$ represents the hazard ratio (HR) comparing the experimental treatment to the control.
We specify the baseline hazard function $h_{0}(t|\theta,\kappa)=(ln2))\theta^{-\kappa}\kappa t^{\kappa-1}$, corresponding to a Weibull distribution parameterized by the median $\theta$ and shape parameter $\kappa$. The resulting probability density function and survival function are $f(t|\theta,\kappa,\beta)=(ln2)\kappa\theta^{-\kappa}{t}^{\kappa-1}\mbox{exp}(\beta Z)$ $\mbox{exp}(-(ln2)\theta^{-\kappa}\mbox{exp}(\beta Z){t}^{\kappa}),$ and $S(t|\theta,\kappa,\beta)=\mbox{exp}\big(-ln(2)\theta^{-\kappa}\mbox{exp}(\beta Z)t^{\kappa}\big).$ The likelihood for patient $i$ with data $(t_{i},\zeta_{i})$ is $L\big((t_{i},\zeta_{i})|\theta,\kappa,\beta\big)=f(t_{i}|\theta,\kappa,\beta)^{\zeta_{i}}S(t_{i}|\theta,\kappa,\beta)^{1-\zeta_{i}}$. Given interim data ${\cal D}_{n}=\big((t_{1},\zeta_{1}),\cdots,(t_{n},\zeta_{n})\big)$ from the first $n$ patients, the joint likelihood function is $L({\cal D}_{n}|\theta,\kappa,\beta)=\prod_{i=1}^{n}L\big((t_{i},\zeta_{i})|\theta,\kappa,\beta\big)$. We assume user-specified priors $\pi(\theta)=\mbox{Gamma}(\phi,\eta)$, $\pi(\kappa)=\mbox{Gamma}(\lambda,\gamma)$, and $\pi(\beta)=\mbox{Normal}(\xi,\tau^{2})$ for the parameters $\theta$, $\kappa$, and $\beta$, respectively. The joint posterior distribution of $(\theta,\kappa,\beta)$ under the user-specified prior is $\pi(\theta,\kappa,\beta|{\cal D}_{n})\propto\pi(\theta)\pi(\kappa)\pi(\beta)L({\cal D}_{n}|\theta,\kappa,\beta)$, which can be sampled using the Gibbs sampler.